Jefferson Lab
JLab Web Server 
Jefferson Lab Home Search Contact JLab

FEL Upgrade Project
September 2002 Monthly Report

Commissioning has started!

We reached a major milestone in the FEL Upgrade Project.  Monday, September 30, is the end date for both the
10 kW IR Upgrade construction project funded by ONR and the first phase of the 1 kW UV Upgrade project
funded by AFRL.  We are continuing a transition from Upgrade construction activities to commissioning activities
that actually began in July.  Discrete goals which the project team had for Sept. 30 included: (1) linac under
vacuum; (2) upgraded injector RF system (100 kW) delivering high power RF; and (3) upgraded high voltage
power supply for the gun (HVPS-600 kV @10 mA) delivering high voltage.  More than 80% of the linac is under
vacuum with the two linac cryomodules and injector cryounit cooled to 2k and ready for start of final SRF
commissioning.  The Personnel Safety System (PSS) was recertified which put all power systems back under
configuration control and allows commissioning activities to proceed.  The injector RF system has already delivered
more than 40 kW to waveguides shorted at the cryounit input.  The interlocks and control power of the gun HVPS
were successfully tested yesterday and today we achieved 100kV.

We hosted several visits of AES engineers to work over design issues on the AES 750 MHz injector cavities.
Substantial progress was made in moving the system engineering detailing forward toward finalization.

We hosted a visit for a group from Boeing to discuss the evolution of our FEL program.  They have long-term
interest in being systems integrator for the Navy drawing on their experience with both FELs and other laser
programs.  We also identified a few areas of potential application interest.

Several members of the scientific and engineering staff attended the International FEL Conference held at Argonne
National Labs this week.  Papers that were presented by the teams and our collaborators at AES, Inc and the
Naval post Graduate School will be added to the FEL bibliography contained in the FEL website: and the JLab Publication website:

We thank the Office of Naval Research for their transfer of the last component of FY02 funds ($350k) for support
of FEL Upgrade activities.

The paperwork was prepared for the required DOC export control declarations on the 1 kW Demo FEL wiggler,
which is to be loaned to the Daresbury Laboratory (UK) for their planned 4GLS advanced light source project.
Copies of this paperwork will be forwarded to the ONR Program Office. The previously prepared loan agreement
has already been sent to ONR.

H. F. Dylla and R. Rimmer from the SRF group participated in a review of the Duke FEL program.  M. Shinn
attended the Boulder Damage Conference and presented a paper on our FEL activities.  G. Neil presented a paper
on THz production at the IRMMW conference in San Diego.

We also verified that our present EA approval covers operation of the IR machine at design levels as well as the
UV machine.  An updated Facility Safety Assessment Document with new safety and operation envelopes for the
Upgrade machine was prepared..  Accel. Div. ES&H and FEL Project Management signed off on a decision
memorandum to extend FEL operations to 30 kW of optical output within the context of the present safety
envelope for the CEBAF site Environmental Assessment.  Copies were sent to the DOE Site Office.

FEL Readiness Process (FRP) for the 10 kW Upgrade is focusing on three EH&S systems: Personnel, Laser, and
RadCon.  Per the previous FEL Readiness Plan of 1998, other aspects of the Upgrade are being assessed for the
Readiness Process by Jefferson Lab and the DOE Site Office.  Findings of the FRP are addressed according to
Jefferson Lab readiness policy: “Readiness Issues” must be resolved before proceeding to next phase and
“Readiness Concerns” are added to maintenance schedules.  The FRP and results will be presented to the FEL
Review Committee in November.

Project Cost Performance:
The Phase 2 FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to
September 30, 2002.  Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k.

Work for Phase 2 proceeded to plan during this reporting period.  Actual charges of $119k were accrued for
September for total accrued costs of $4,202k to date for Phase 2.  Work scheduled for September was $505k for
a total of $4,442k for Phase 2 to date; work performed for September was $390k for a total of $4,184k for
Phase 2 to date, which results in 100% scheduled vs. 94% performed for the Phase 2 project.  At present the cost
variance for Phase 2 is -$19k and the schedule variance is -$258k.  We have met the planned cost at completion
of Phase 2 effort.  All but 6% of the scheduled work (earned value) has been booked at the completion date.
Remaining items in the original project scope have not impeded the start-up of the FEL Upgrade commissioning
which began in September. We are actively working with the vendors who are supplying the remaining hardware so
that the delivery and integration of this hardware can be interleaved with the commissioning activities over the next 3
months and not impede the commissioning schedule.

The UV FEL project (Phase 1) was approved for $2,836k for a performance period from September 2001 to
September 30, 2002.  Actual charges of $152k were accrued through September for a total accrued cost of
$2,221k since project start.  Work scheduled for September was $458k for a total of $2,836k for the UV project
to date, work performed for September was $550k for a total of $2,568k for the UV project to date, which
results in 100% scheduled vs. 91% performed for the UV project.  At present the cost variance for the UV project
(Phase 1) is $347k and the schedule variance is -$269k.  The current
cost and schedule variances will clear when the last quarter of charges for the Aerospace subcontract are costed in
December due to the 3 month late start on subcontract.

Phase 2 of the UV FEL Project was approved for  $1,472k for a performance period from August 2, 2002 to
June 30, 2003.  The start date was 4 months later than planned due to late arrival of AF funds.  Actual charges of
$697k were accrued through September for a total accrued cost of $1,078k since project start.  Work scheduled
for September was $8k for a total of $16k for the UV (Phase 2) project to date, work performed for September
was $460k for a total of $678k for the UV (Phase 2) project to date, which results in 1% scheduled vs. 46%
performed for the UV (Phase 2) project.  At present the cost variance for the UV (Phase 2) project is -$400k and
the schedule variance is $662k.  Because of the late arrival of funds significant expenditures had to made early in
the first 2 months which have generated both the current cost and schedule variances.  Funds for this project are
essentially fully obligated and we expect to track remaining earned value on the UV optical cavities and the Phase 2
Aerospace subcontract which is scheduled to be costed in the next quarter.

WBS 4 (Injector):
All the hardware for the ball cathode has been silver-plated and about 70% of its components are ready for
assembly.  Some rust spots were found on the shield door mechanism during the cleaning process.  It was also
found that this piece became slightly magnetic.  A new piece is being fabricated with 316L stainless steel.  The light
box has been partially cleaned and all its internal components silver-plated.  The resistor drawings have been
submitted to the machine shop. Polishing of the high voltage ring is in progress.  The stalk tube is being smoothed
out with sandpaper in preparation for vacuum.

Field emission scans were performed on two stainless steel one-inch discs polished down to one-micron finish.
The results show ten emitters in the range of 80 to 140 MV/m for the bare stainless steel sample, while only two
emitters were found around 120 MV/m on the ion-implanted sample.  The maximum field in the cathode support
tube, which has already been implanted will be around 12 MV/m.

Gun HVPS - The large toroid was aligned, the resistor mounting bracket made and welded onto the toroid by the
Machine Shop, and the toroid reinstalled back in the tank this week.  A system diagram is about half done for the
wiring and the interface to the PSS & MPS.  The HVPS for the Gun was then assembled and passed PSS
Certification.  It was successfully operated at 100 kV in air.  An additional SF6 bag will be purchased to contain
the additional volume of the new tank.

The rear bellows/drive unit and rear corona shield were submitted to the shop for fabrication. The gun stand was
grouted in place and set up / cleaning of the clean enclosure around the gun chamber was done.  Assembly of the
electrode structure into the ceramic stack was started in the off-line clean room of the test lab.  Modifications were
made to the gun stand to correct an interference problem.  The ground end internal corona ring is being polished
and nearly complete.  Assembly of the light box has been started.  Assembly revealed some offsets which are being
corrected by re-machining of parts.  Field measurements of the guns reworked solenoid were started.

We developed a problem with the gun assembly and have incorporated a work around with minimal impact on the
commissioning schedule.  During the assembly of the upgrade gun cathode structure, we noted an exfoliation of the
high voltage conditioning film that is deposited onto the cathode ball.  Previously, no deterioration has been
observed with this film on stainless steel substrates including on a large (100cm2) test sample that was coated 3
years ago and stored in ambient conditions.  After some analysis of our records and procedures, it was determined
that the localized coating failure was due to the use of an Al support rod which supported the ball during
deposition.  The film adhesion on the opposite side of the ball was unaffected.  The action plan that was initiated
immediately after the noted failure was the following:
(1) the film was removed from the cathode ball using a stainless steel etch that was previously tested when we had
to remove trial films from the cathode support tube in July and August.
(2) we began re-polishing the ball in preparation for a re-deposition
(3) the deposition chamber was readied for use including machining of a stainless steel support in place of the
original aluminum support.

It is estimated that the ball preparation procedure will take a week followed by another week for the 2-step
deposition process.  The process chamber has recently been upgraded with a larger vacuum pump and we note
that the geometry of the ball allows this deposition to be done relatively quickly compared to the extended
cylindrical geometry of the cathode support tube. This two week delay impacts the final assembly and pump down
of the gun, however the delay should not significantly impact the overall commissioning schedule.

The drive laser transport line was completed up to the last mirror can and pumped down.  Installation of the final
tube with the exit Brewster window must wait for light box installation.  We mounted the ion pump on the light box
and installed the first of four mirrors in the box.

WBS 5 (SRF):
Construction of the cavities for the third FEL cryomodule is fully underway:
    - Cavities JL010, 11 are welded into seven cell structures
    - JL012, 13 are being prep’ed for dumbbell welding
    - JL14,15,16,17 cells are undergoing inspection and polishing

We summarized the measurements of EP3 in the vertical test and of the copper model with additional elliptical
HOM probe on the FPC end.  The FPC termination seems to produce a useful improvement in one orientation of
the TE111#5 mode at 1871 MHz but otherwise no major differences with previous HOM-only data.  This should
at least raise the threshold of one mode in the FEL.  The elliptical HOM probe added to the FPC end of the
copper model also produced a useful improvement in one bad mode, #11 in Kevin's table, raising the threshold by
almost a factor of 10, and improved mode #13 by about a factor of two.  Unfortunately it had no real effect on
mode #5, at least in these measurements. In some cases the Q's are higher than EP3 despite the added HOM
antenna, which may be due to the lack of field flatness of the copper model.  Inspection of the SL21 string shows
how little room there is to make any changes to the end groups.

Given the results to date and the technical difficulty of moving the couplers closer and the potential schedule impact
of delays trying to improve damping further we have decided to proceed with construction of the cavity as presently
designed.  We will continue to study the damping for possible improvements in future modules.

WBS 6 (RF):
Quarter HVPS - All of the RF couplers have been calibrated and reinstalled.  The LCW manifolds for the
Circulator, RF Load, and HOM filter have been plumbed and leak tested.  Unit 3 was turned On under local
control and ran 44 kW of RF into a short at reduced cathode voltage.  The voltage taps were then set to 95% of
full voltage and achieved 34 kV with the klystron load of 6.4 amps.  This voltage is more than enough to achieve
110 kW for the klystrons.  The ripple voltage is 510 volts p-p at full power.  Unit 3 also passed the PSS
Certification.  Unit 4 HVPS and klystron cart are still being assembled and will be tested within the next 2 weeks.
The system prints are being updated and nearly ready for release.

Buncher - The Buncher water skid controller is not connected to the Buncher cavity or the LCW.  This is to relieve
congestion for the installation of the Gun and Light Box.  It passed the PSS Certification with a jumper for the
LCW interlock on the water skid.

Zone 1 - The waveguides were reconnected to the Quarter.

Zone 4 - The half height waveguides were completed by the Machine Shop and brought over to the FEL.  They
will be installed next.  Due to a bad connection on the waveguide interlock, this zone did not pass the initial PSS
Certification test.

Zone 2 - The waveguides for this zone were reconnected again this week.  It passed the PSS Certification.

WBS 8 (Instrumentation):
The mechanical design for the recirculation dump (4") Beam Position Monitor (BPM) is complete and passed on to
the mechanical designers.  This new design will feature a ISI Seal SMA feedthrough on a mini-conflat instead of the
welded feedthroughs that were prone to failure during assembly.

The design for the Personal Safety System (PSS) interface to the new gun high voltage power supply is complete.
The new interface chassis is progressing well, this new design will greatly simplify the operation and trouble shooting
of any high voltage problems.

The Linux frame grabber software is progressing nicely. Further work has been done in writing code to find the
reference marks, and measuring the spot size.

The schematic for the Acromag fanout module has been given to the Fast Electronics (Physics - THANKS) group
for layout.  This will be used both for high voltage power supply control and for the drive laser control.  The
checkplot and schematic for the isolation module fan out board have been received from EECAD. Cables were
pulled from the stepper motors chassis rack to the Optical Cavity and Lab 2.  The software group has started
updating the magnet controls for the 0F region using the RS-485 communications upgrade.  The schematic and
plots for the Acromag fanout module have been completed, reviewed and awaiting sign-off.

We are continuing investigating interest for commercial engineering firms to bid the regulator/DSP board design and
assembly for a new trim power supply.  There are a number of firms that have excess capacity since the down-turn
of the telecom market and are hungry.  We are pursuing purchasing a 60 amp current transducer from Danfisik as
an OEM part.  They are producing ~200/month for MRI shim magnet power supply regulators.  This device would
have 50 - 100X better performance (stability/reproducability) than the LEM module used in the Fermi switcher at a
lower cost.

Progress has continued for the design of the HV interlock chassis and fabrication has begun.  The HV interlock
board is near completion.  Cable tray was hung above the SF6 tank and cabling for the HV has begun.

Plots for the GPIM fanout board have reviewed and ready for sign off. Fabrication of the Stepper Motor breakout
chassis has begun.  Two (2) Beamviewer Canbus Interface modules have completed and turned over to A. Grippo
for testing.

BLMs - The F0150 rev A (rear cards) back..  All parts are in house and ready for assembly.  The F0150 pcbs
(front cards) are proceeding well thru the testing phase. The F0150 rev A (rear cards) are back from the board
house and look great. We immediately assembled one to serve for the first 12 BLM channels to come online by
Oct. 1st. The remaining boards will be implemented later in the commissioning process as priority demands.  The
F0151 rev - prototype is fully populated and mostly programmed.  The card set is currently installed into and
communicating with iocfelt2 in Lab 5 where it will finish out its testing next week and then be installed into the FEL
proper. Testing of the BLM prototypes is almost complete.  The cards are being tested with the actual BLM heads
and the LeCroy high-voltage power supply.  They will next be calibrated and then installed.  This will provide the
baseline functionality for commissioning for 12 channels, after which a second set will be built-up and further
developed to provide the advanced features such as the "beam veto during the macropulse" and the "trip
sequencing" functions that are not fully developed.  Once the firmware for these function has been developed, the
1st set will be 'flashed' with the upgrade.

DLPC - The CO305 carrier board pcb layout and schematic are finished and the package has been sent out for
production on a 5 day turn.

MPS - PCB layout and schematics were checked and revised for the 6U breakout adapter (for the general
purpose isolation module).  The new MPS boards have gone out for fabrication.  These will eliminate the need for
any masking of hardware devices in the upgrade.  We began the MPS rack rework, this eliminates the patch
cables, PLCs, and terminal blocks.  The design was completed for the control room interface box.

MISC - The photocathode scanner chassis is being recommissioned in Lab 5. Assistance in documenting the high
voltage power supply design was provided as well.

The vacuum rack is nearing completion, currently the ion pump power supplies are downstairs local to the pumps.
Next week the 3 racks will be recommissioned as well as the new software and all the controls and readout will be
available in EPICS.

The dump water skid was refilled and re-commissioned.  One of the conductivity monitors had failed and is being
replaced, both of the sensors were replace as preventative maintenance.  Railings and crossovers were installed in
the 3F region to prevent any trip accidents for people trying to get over the beamline.

The interlocks to the gun high voltage power supply are now complete.  The safety group did the PSS certification
and we have begun to turn on the various systems.  AC power to FL01B08 and FL05B07 and 08 has been
completed with the breaker boxes upgraded to support 208, 3 phase power.  These racks support the Injection
and Extraction dipoles.  The  Injection and Extraction 5 kW Danfysiks power supplies also were installed in Zones
1 and 5 respectively.  Water hookups to the PS have been completed.  Shunt chassis for the First Arc String have
been installed in FL05B09 and the wiring is in progress.  Radcon has begun work in the vault for the Rapid Access
system.  Four GP 64 bit VME Isolation modules are being populated. Assembly has begun in the 0f and 1F regions
for the Beamviewer camera assemblies.

Parts have been ordered for 40 additional beamviewer control boxes.  And we are in the process of stuffing more
of the GPIM64 VME boards for installation in various systems.  EH&S staff and Deborah did a safety
walkthrough, no major deficiencies were noted, however general housekeeping could use some improvement.

We began a collaboration with MIT Bates lab to get EPICS working on Linux.  This has been done by many
groups using the X86 chips but our focus will be with the new Motorola Power PC chip MPC8266ADS-PCI.  In
fact our group is using a Dell PC and framegrabber as an IOC to process the viewer video.  We worked together
to establish a clear path that will result in a low cost - license free IOC.  This PPC chip has PCI bus support
allowing the use of countless commercial I/O boards.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
•  The measurements of the bulk field on the first GV continued, finally getting repeatability after
    going between energies in the part in 10^5 level.  Field flatness is within 2 parts in 10^4.  An
    error in overdriving the magnet core lead us to use the trim coils to erase the magnetic memory of
    the magnet in one stroke rather than go through a 5 hour degaussing procedure.  That led us to
    using a similar procedure between all magnet settings.  It works beautifully.
•  By using various shims on the backside of the field clamp edges on either side of the center, we
    were able to bring the integral linearity (It had sextupole) of the GV to within 2 parts in 10,000.
    (Observed with a course scan.)  When we used the fine scan, the integral linearity is 4 parts in
    10,000 and we can see that we have a very powerful tool to adjust the field in to specification
    with the shims.
•  The first GV qualified for the machine and was signed off by Dave Douglas after two additional
    iterative magnet measurement learning cycles.
•  We had solved the linearity of the field integral as you measure in transverse positions (integral
    gradient) last week.  We homed in on the values of that field integral in the transverse positions,
    obtaining the correct values to within 2 parts in 10,000.  In doing so, we calibrated our spread
    sheet calculations to dial the micrometric field clamp adjustment values.
•  Formal measurements to this point explored and finally obtained the nominal 10 MeV/c
    requirements of 555.940 Gauss and a half sagitta chord integral of 11673.9 g-cm.  When we
    went to explore the 9.5 MeV/c bulk field setting the field integral didn't scale by 1.8%- way out of
    Dave's specs.  The dynamic range of these magnets is very small because they depend on
    hysteresis in the yoke.  As a result, David Douglas redefined the specification for this magnet style
    to operate at 10,750 g-cm, (9.2 MeV/c) the value the IR Demo successfully operated at.  We
    quickly homed in on this integral and the corresponding bulk field (by shifting the magnetically
    defined centerline of this wedge magnet) and obtained bulk field flatness and integral gradient to
    within Dave's requirements and signed off the magnet as qualified.
Optical Chicane Dipoles (GW)
MEI discovered the reason behind the gap height problems with the second batch of cores and is having the return
legs touched up at the grinders.  The coils are finished.  We await the successful core check out and glue down of
the Purcell gap materials and we will have a good
magnet batch.  Magnet Test will activate the field integral match stand on two GW magnets  to start their formal
Arc 180 Degree Dipoles (GY)
•  GY Coil Potting fixture continues its preparation for potting.
•  We resolved our contractual issues with Wang NMR over winding the replacement for the first
• We were notified of the shipment of the extra copper for the fifth coil from Finland.
•  The extra copper for the third and forth layers of the 5th coil was received by Wang NMR.  They
    had completed winding the first two layers earlier in the week.
•  Wang started preliminary work on gluing the shims down on the GY cores.
•  Wang NMR continued winding the remaining two layers of the last GY coil.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
•  The GX cores were shipped  to Wang NMR. They are in final assembly.  One glitch in the cover
    bolt pattern became evident and is being fixed by PECo.  We have had very few glitches of this
    sort, a tribute to the quality of the drafting and the use of relational 3 dimensional drafting models.
• At Wang NMR, the first GQ coil was vacuum impregnated with epoxy.
•  At Wang NMR, the second GQ coil was vacuum impregnated with epoxy and first coil passed
    hi-pot test.  They started assembly of the GQ magnet.
Trim Quad (QT)
•  We started formal measurement of four trim quads on the test station after resolving the problem
    with the test stand.  The temperature problem with the side coils has been fixed with the new
    cooling plates.  All four trim quads were fully measured.
Sextupole (SF)
• DULY Research didn't finish the drawings as they had indicated they would.  We pulled them as is
    and will deal with them in some other way.  DULY Research passed all the magnet drawings to
• We have decided to fore-go immediate fabrication of the SF sextupoles for eight locations in favor
    of using modified and doubled-up IR Demo Sextupoles in two locations in the 2nd Arc.  David
    Douglas believes this configuration will still allow lasing, but at a small reduced intensity during a
    commissioning period.  This grace period will allow time for fabrication of the SF sextupoles.
    They can then be installed later to gain full intensity.  We are working on the design of the quick
    fix sextupoles.  A preliminary layout indicates that the quick fix sextupoles can be made from the
    IR Demo sextupoles and the simple addition of some shim pieces, some altered coil supports and
    some redrilling of the original core pieces.  We passed a preliminary layout of the quick fix
    sextupoles to Process equipment, the folks who had made the original IR Demo sextupole cores
    for comment on the shimming.
Octupole (OT)
•  The Octupole design is about half complete at DULY Research while they finish the Sextupole.
Beam Line and Vacuum
•  X-Chamber - The third transition (with small machining gouge) was approved for use in its
    chamber.  Final rework and machining on the x-chamber and transitions (for weld preparations)
    is in progress.  Welding of transitions to chamber progressed.  Fabrication of the chamber is
    finished and it is being cleaned and leak tested by the Vacuum Group.  Welding of transitions to
    chambers was a challenge.  We needed to remove material in order to obtain a full penetration
    joint.  The joints were re-welded.  The experience learned from this effort will be used to modify
    the weld prep on the other chamber transition interfaces.
•  Chicane Chamber - Wire EDM work for the transition piece has been quoted.  Chamber plates,
    tubes, and flanges are on hold to allow higher priority parts to be fabricated. Wire EDM work for
    the transition piece has been awarded.  Welding of the first chicane chamber was started.
•  Y-Chamber - All material for the Y chamber has been received.  Plate machining is in progress
    and is almost complete.
•  ARC Chambers - Master Machine put together and welded the first chamber.  All the flanges
    have been machined.  The last transition is on the wire EDM machine. Master Machine has
    completed the fit up of the first chamber.  For the other ARC chambers, the machining of the
    bevels was delayed due to a controller problem on the CNC machine. The machine has been
    repaired and is being run through some initial testing.  A subcontractor has been utilized to keep
    machining efforts moving.  An alternate configuration for the paddle viewer ports was approved
    and will be implemented into chambers using redline updates of drawing dimensions.
•  The stands for sextupoles and trim quads were received.
•  We signed off the location drawing for the extraction stands.
•  Low Conductivity Water piping installation continued in the FEL enclosure, it is now more than
    60% complete.

WBS 11 (Optics):
Mechanical work on the optical cavity was spent primarily measuring the mechanical resonances of the 3"
deformable mirror assembly.  This information is being used to better model the feedback circuit required for the
active mirror steering.  We found a commercial coating vendor to sputter tungsten disulfide onto our leadscrews.
Two leadscrews are being stripped of their silver coating and will be sent to the vendor for coating.  The 2"
outcoupler mirror mounts (2") arrived over two weeks early and are being prepared for the subsequent brazing and
plating steps.

Detailing continues at a rapid rate on the optical transport turning cassette, the collimator and insertable mirror
designs.  As drawings for subassemblies are complete they are checked and sent out for procurement.  We are
putting together orders for other parts that will be used for these assemblies.  Optical modeling shows that the
additional distance between the outcoupler and the collimator does not necessitate a complete redesign of the

We assisted in reinstallation of optical components for the drive laser optical transport.  We received spare mirrors
for the transport, and they will be QA'ed.

Water cooling was supplied to the 3" optic in the test and assembly stand, and the amount of noise quantified.  The
mass of the yaw plate effectively dampened any noise on that axis.  For the pitch plane, the noise in the ~ 70 Hz
region increased by ~ 15 dB.  When driven, the resonance frequencies shifted some as well, as was expected.
Given the small amount of additional noise, the feedback circuit is being modified to control it.  The design of the
turning cassettes has made great progress, and we have ordered a large number of parts needed.  This was also
true for the insertable mirror.

The leadscrews for the optical cavities have been sent to Thin Films, Inc, to be sputtered with tungsten disulfide.
We received all the LVDTS, prepared for high vacuum use, required for the optical cavity and optical transport to
the OCR.  Outgassing checks are ongoing at the time of this report.  The OCMMS feedback control is
progressing, with dampening by about a factor of 2 in the 1-15 Hz region, in a very noisy environment.  Parts of the
optical transport hardware are being procured.
We installed a HeNe in the drive laser clean room in order to align its optical transport.  The transport is aligned to
the last turning mirror over the light box.  Some work remains to insure everything is firmly attached before the relay
telescope is installed.  The optical components for the photocathode scanner were modified for the new light box
and are ready for installation.

The tungsten disulfide sputtered leadscrews for the optical cavities were tested under load this week.  Initial tests
looked quite good, with a measurement of drag torque of ~ 20 in-lbs.  However, after a few iterations, the drag
torque increased.  Our initial impression, based on microscopic inspection, suggests the film is wearing off.  While
we will continue to investigate this, for now, we intend to move forward with the use of a Vespel nut/bare stainless
steel leadscrew combination.  The LVDTs have been checked for outgassing, and are sufficiently good for the IR
optical cavity and transport system.  Outgassing checks on other components are ongoing at the time of this report.
We received the HR mirrors for 6 micron operations.  They are being inspected.  A prototype outcoupler mirror
holder was nickel-plated prior to brazing of the cooling tube.  Some chipping of an edge (in a noncritical area) was
noted, and suggests we want to make that edge thicker.  The OCMMS beamline hardware was sent to the
machine shop for modifications to one item, and final UHV cleaning.  When they return, they will be installed in the
beamline.  Procurement of optical transport hardware is proceeding at a rapid rate.

In preparation for commissioning activities, a leak in the drive laser beam dump was repaired and the drive laser
operated.  It is operating at above spec power.  Other specs will be checked next week.  The drive laser optical
transport was aligned from the drive laser enclosure to the last turning mirror above the light box.  At this time the
transport is being pumped out to eliminate water, and will be backfilled with dry nitrogen.

Aerospace/Jefferson Lab Subcontract
We used finite element modeling software tools to evaluate the effect of a XY stepper moving a 50 Kg mass, 0.5m
at velocities of 20 nm/sec.  It was expected that this motion would induce deflection of the SS plate of a vibration
isolated honeycomb table.  The models show that the deflection is on the order of 1 ?m, well within our
specification.  We also calculated the lowest order modes of a 4' x 8' or 5' x 8' optical table.  This was to insure
that we could add appropriate dampers or make sure that we stay away from these table resonances in our design
of the optical infrastructure.

We refined the design of using 2 Pockaell cells for both pulse picking and intensity modulation of a 5W FEL (UV)
beam.  The driver circuits required were analyzed as was the I/O output available from the XYZ stepper motors to
see if we could get dynamic analog/digital output that is directly related to the XYZ vector velocity.

Finally, we analyzed the use of high bandwidth tip/tilt piezo transducers to stabilize a possible vibration of the FEL
Laser beam and or optical mounts.  A lab demo is being setup.

                                  Other Related Projects:

IR Upgrade Commissioning (ONR)
This project was funded in FY02 with $1,000k, with a period of performance from July 1, 2002 to March 30,
2003. Additional funding of approximately $1,000k is anticipated in FY03. In this reporting period, $150.7k has
been expended for a total expenditure of $572.4k.  Commitments total $153k.
See Highlights section above.

Crymodule 3 (ONR)
This project was funded in FY02 with $400k.  Additional funding of $350k is anticipated in FY03.  In this
reporting period, $151.3k has been expended for a total expenditure of $230.7.  Commitments total $30.4k.
See WBS 5 report above.

One micron wiggler project(ONR)
This project was funded in FY02 with $600k.  Completion funding of $600k is anticipated in FY03.  In this
reporting period, $34k has been expended for a total expenditure of $111k.  Commitments total $0.7k.
A poster presentation was given at the 2002 FEL conference in Argonne.  The paper has been accepted for
publication.  While in Argonne, a meeting was arranged between a representative from STI Optronics and
engineers from Jefferson Lab to discuss potential design specifications for both the wiggler and the gap mechanism
for the one micron wiggler.  STI Optronics made a 5.5 cm period wiggler for Argonne and feels that they can
modify the pole design to produce a 6 cm period wiggler for us.  This design has the advantage of not pushing the
field up so high so that less expensive pole material may be used.  We also took advantage of the location to
inspect the new undulator A gap mechanism now in use on the APS.  This design is very simple mechanically but a
potential nightmare from a controls standpoint.  Since APS is committed to making the control system work the
main problem is that the savings in the mechanical design will be offset in increased control system costs.  STI also
has a design that is overkill for the one micro wiggler but that is already designed and has a very simple control
system.   Either design will meet the desired specifications for the one micron wiggler.

JTO Drive Laser Project
After verifying the claimed performance versus vendor claims we are in the process of proceeding on the major
laser hardware purchase.
We await a quote from Time-Bandwidth and are ready to go forward with a purchase request on receipt of that
updated quote. No funds were expended this month while awaiting their response.  We are ready to prepare a sole
source procurement for the Master Oscillator.
During this reporting period expenditures of $4.3k were incurred.  Commitments total $0.0k.  Total expenditures to
date are $23.5k.  The program is on track financially and is expected to deliver all elements of the SOW within
projected cost.
JTO Cryomirror Project
We continued efforts for proceeding with engineering tests of the proposed system.
We received the components for mounting a cryocooled sapphire mirror (brazed into a molybdenum holder) so as
to do temperature-induced stress birefringence tests.  Measurements will be performed into October.

We received a prototype holder for the UV & IR Upgrade optical cavity.  It was Ni-plated and vacuum annealed.
Brazing the cooling tube will occur in October.  We also designed two adapters that will allow us to mount the 2”
dia mirror in the 6” dia gimbal mount that is in the Mirror Test Stand.  We anticipate these holders (one made of
G-11 fiberglass epoxy, the other of Macor glass) in late October.  Measurements will occur in late October or
early November.
During this reporting period expenditures of $47.3k were incurred.  Commitments total $132.1k.  Total
expenditures to date are $106.3k.  The program is on track financially and is expected to deliver all elements of the
SOW within projected cost.